OpenStudy (anonymous):

How do you solve nCr•nCr/nCr? its in a fraction form and i dont understand.... : (6)C(2)•(9)C(1) / (15)C(3)

jimthompson5910 (jim_thompson5910):

6 C 2 , 9 C 1 and 15 C 3 are all integers. For example 6 C 2 = (6!)/(2!(6-2)!) = (6!)/(2!4!) = (6*5*4!)/(2*1*4!) = (6*5)/(2*1) = 30/2 = 15 So 6 C 2 = 15 So simply evaluate each combination expression using the same idea shown above and simplify as it were a normal expression.